# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 3.3.2 CrowS-Pairs
Dataset Taxonomy. CrowS-Pairs adopts a contrastive sentence-pair structure, where the only substantive difference between each pair is the demographic group mentioned. Each pair is labeled as “stereotypical” or “anti-stereotypical” depending on whether the first sentence reinforces or challenges a stereotype. The dataset was crowd-authored via Amazon Mechanical Turk, where workers were provided with partial prompts (e.g., from MultiNLI or ROCStories) and instructed to compose minimal pairs reflecting societal biases. Substitutions were guided by a predefined list of group pairs (e.g., “Black” vs. “White,” “women” vs. “men”), and all submissions underwent a rigorous validation process involving five annotators. Only examples with consensus on minimal edit distance and stereotype direction were retained.
The dataset is monolingual (American English), stylistically diverse (including narrative, conversational, and factual tones), and grounded in naturalistic contexts. Sentence construction aimed for minimal edits outside of group substitutions, though some stereotypical sentences may appear more fluent—a reflection of societal linguistic priors. CrowS-Pairs targets nine demographic bias categories: race/color, gender/gender identity, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each pair is tagged with a primary bias type.
@@
Representativeness Bias. Following the methodology defined in Section 2.3, we assess representativeness bias in CrowS-Pairs by comparing its demographic distribution to U.S. population priors from the 2024 Bureau of Labor Statistics. Let $P_{\mathcal{D}}(A)$ denote the empirical distribution across demographic categories and $P_{\mathcal{P}}(A)$ denote the reference distribution. The representativeness bias is computed as Eq. 1, the Kullback–Leibler divergence between these two distributions. This metric reflects the extent to which the dataset over- or under-represents certain demographic groups relative to real-world proportions.
Race–Color. CrowS–Pairs significantly over-samples sentences targeting Black individuals (40.1% vs. 13.6%), while under-sampling White (44.3% vs. 58.9%) and Hispanic groups (5.3% vs. 19.1%). This results in a high divergence of
@@
Gender. The gender axis is more balanced, with a slight male overrepresentation (55.1% vs. 51.2%) and female underrepresentation (44.9% vs. 48.8%), yielding
@@
Other Attributes. While CrowS–Pairs lacks structured subgroup labels for attributes like socioeconomic status, nationality, disability, and religion, visual comparison with the BLS reference distributions (see Figure 9) reveals similar mismatches:
@@
Figure 9: CrowS–Pairs (blue) vs. 2024 BLS demographic distributions (red) across nine protected attributes. Over- and under-representation are evident where bars differ substantially.
Annotation Bias. CrowS–Pairs relies on crowdworkers to generate and label sentence pairs as either stereo (stereotype) or anti_stereo (counter-stereotype). While validation included consensus filtering from five annotators per item, the high rate of stereotypical labels across most categories suggests potential annotation bias.
We define annotation bias as the maximum difference in stereotypical annotation probabilities across bias types or demographic groups (as in Eq. 2). A detailed audit over all 1,508 validated pairs yields the following findings:
  • Bias Type Level: The stereotypical annotation rate varies substantially across bias types. Disability (0.9500), religion (0.9429), and race-color (0.9167) exhibit the highest rates, while gender (0.6069) and physical appearance (0.8254) show notably lower rates. The maximum inter-type annotation gap is:
@@
    observed between the gender and disability categories.
  • Demographic Level: Group-specific annotation bias also appears within protected attributes. For race-color:
@@
    The maximum intra-attribute gap occurs between Asian and Other:
    $B_{\mathrm{ann}}^{\text{group}}=\max_{g_{i},g_{j}\in\text{race-color}}\left|\Pr[y=1\mid g_{i}]-\Pr[y=1\mid g_{j}]\right|=0.0816.$ (26)
    Within the gender attribute, stereotypical annotation rates are 0.8440 for male and 0.8031 for female references.
These results suggest that certain categories—especially disability, religion, and race-color—are more prone to being annotated as stereotypical, potentially reflecting both annotator bias and societal salience of these stereotypes. By contrast, gender shows relatively balanced annotations and the lowest $B_{\mathrm{ann}}^{\text{type}}$, making it a potential control category.
The high stereotypical labeling rates across nearly all demographic groups indicate a structural inclination among annotators to reinforce stereotype-based assumptions. Such annotation bias may affect downstream model bias evaluations and should be accounted for in fairness-sensitive applications.
Stereotype Leakage. We computed pointwise mutual information (PMI) for every (group, trait) co-occurrence using a fixed vocabulary of protected-group and trait tokens. The word lists span five protected categories (race–color, gender, religion, age, and disability) and four trait domains (occupation, personality, social_status, and behavior). Figure 10 visualizes the results.
@@
  • Mutual-information ranking (bits): gender–personality ($1.04$) $>$ gender–behavior ($0.92$) $>$ race/social_status ($0.54$) $>$ religion/occupation ($0.50$); combinations such as race–behavior and age–social_status yield $\mathrm{MI}=0$, reflecting either scarcity or filtering.
@@
Differential Metric Bias. Following Eq. (8), we computed the paired score gap $\Delta_{f}=x^{(\mathrm{more})}-x^{(\mathrm{less})}$ for seven auxiliary metrics over all $N=1{,}508$ sentence pairs. Figure 11 portrays the mean gap for every bias category.
[…]
Most Perspective-API facets (toxicity, severe_toxicity, identity_attack, insult, threat), VADER sentiment, and Regard show no statistically significant overall gap after Bonferroni correction ($p>0.05/7$).Toxicity: $\bar{\Delta}=0.021$, $t=1.37$, $p=0.18$; Sentiment: $\bar{\Delta}=-0.039$, $t=0.04$, $p=0.97$; Regard: $\bar{\Delta}=0.0004$, $t=-1.49$, $p=0.14$. The lone exception is the gender-polarity suite, where all three variants yield small but reliable negative gaps (unigram $d=-0.26$, max $d=-0.34$, wavg $d=-0.28$; all $p<5{\times}10^{-4}$). Stereotype sentences therefore skew slightly masculine in embedding space, despite containing no overt gender terms beyond the protected token swap.
Figure 11 reveals that metric bias is heterogeneous across bias types:
  • Gender rows illustrate the massive negative gender_unigram gap ($-0.261$), corroborating the aggregate significance of gender polarity.
Because most classical toxicity metrics appear unbiased in the aggregate yet display sizeable, topic-dependent deviations (Fig. 11), we caution that per-category auditing is indispensable. The reliable yet small masculine shift in gender polarity confirms that embedding-based metrics themselves encode residual gender stereotypes (despite content control), echoing the instrument bias concerns of Section 3.5. Model fairness studies using CrowS-Pairs should therefore: 1) report both global and category-level $\Delta_{f}$, 2) disclose auxiliary-metric validation, 3) and interpret gender-related findings in light of the metric’s own bias.
@@ 3.3.3 RedditBias
Dataset Taxonomy. Structurally, RedditBias offers two levels of granularity for evaluation: full Reddit comments (up to 150 characters) and phrase-level windows (±7 tokens around a target term). These units serve as prompts to probe model responses. The dataset was compiled using 3.33 years of Reddit data obtained via the Pushshift API. It targets four demographic axes—Gender, Race, Religion, and Queerness—using curated lists of group identifiers (e.g., “women,” “Muslims”) and negative attributes (e.g., “manipulative,” “mentally ill”). Phrases such as “gays are mentally ill” were retrieved using pattern-matching queries. Extracted data was filtered for readability and toxicity, with bias expressions centered in the phrase windows. The language is informal, often offensive, and rich in sarcasm and slurs, reflecting the authentic linguistic style of Reddit. This real-world grounding distinguishes RedditBias from crowd-sourced or templated datasets. The dataset focuses on overt, toxic stereotypes and avoids subtle or positive bias expressions. It is distributed under a CC BY-NC-SA license and includes verbatim Reddit content, which may limit certain applications.
@@
Strengths and Limitations. A key strength of RedditBias lies in its realism—it reflects how bias and hate speech appear in natural online dialogue. The dual-level design (comment and phrase) enables both coarse and fine-grained assessments. It also supports both intrinsic scoring using model-likelihood metrics and downstream evaluations of dialogue performance. However, the dataset has several limitations. Its domain specificity may reduce generalizability to other platforms or languages. It focuses solely on four demographic dimensions and explicitly targets harmful bias, omitting subtle or structural forms. The reliance on pattern-matching during data collection may overlook implicit stereotypes, and the use of sarcasm or ambiguous phrasing complicates annotation and model interpretation. Additionally, the dataset contains highly offensive content, which necessitates ethical caution in use and publication. Its composition is heavily skewed toward toxic expressions, without balancing benign or neutral dialogue, which may influence evaluation metrics.
Bias Analysis. RedditBias presents a unique case for bias analysis due to its real-world, unfiltered origin. It is particularly suited for evaluating a model’s exposure to and propagation of harmful stereotypes in dialogue generation. Specific metrics such as model-generated likelihood scores, differential response tendencies, or sentiment shifts across demographic prompts can be employed to quantify the bias exhibited during generation. Further analysis following the framework in Section 2.3 can reveal representational imbalances or model vulnerabilities specific to conversational settings.
Representativeness Bias. Using 2024 U.S. Bureau of Labor Statistics (BLS) benchmarks we recomputed the KL divergence between the RedditBias attribute distribution and the reference population, obtaining
$B_{\text{rep}}^{\text{RedditBias}}=D_{\mathrm{KL}}\!\bigl{(}P_{\mathcal{D}}(A)\,\|\,P_{\text{BLS}}(A)\bigr{)}=17.75,$ (27)
far larger than both our earlier estimate (14.51) and the Wikipedia-based baseline reported in this section. The principal drivers of this divergence are:
  • Gender. Female-coded terms (girl, mom, wife, lady) dominate, whereas BLS reports a near-parity labour distribution (51.2 % male, 48.8 % female).
  • Sexual Orientation. LGBTQ + terms (gay, transgender, lesbian) appear in $\approx$ 7× the proportion suggested by BLS figures (7.2 % LGBTQ +, 92.8 % straight).
  • Race / Ethnicity. Black-related tokens are heavily over-sampled; Whites (58.9 %) and Hispanics (19.1 %) are markedly under-represented relative to BLS population shares.
  • Religion. Jewish and Islamic terms occur much more often than their combined 37 % share of non-Christian affiliations (BLS: 63 % Christian, 37 % non-Christian).
@@
Annotation Bias. We re-evaluated annotation bias with the automatic label $g_{\theta}(x)=\mathrm{ppl}(x)$, the mean sentence–level perplexity produced by a 6-layer GPT-2 LM fine-tuned on balanced news.Perplexity serves as a proxy label in many toxicity pipelines because higher values correlate with out-of-domain or stylistically deviant content. Using Eq. (2), the maximum groupwise gap is
@@
far exceeding the 95% bootstrap confidence band for neutral vocabulary ($\approx\!1200$). Detailed findings are:
  • Gender. In the gender files the largest gap (23,675.67) occurs between niece and bride; class means vary from 803.4 (niece) to 8,424.4 (mother), indicating the LM treats female-coded tokens as markedly harder to predict.
  • Sexual Orientation. A gap of 3,622.6 separates homosexuals and bisexuals, with lesbian scoring 4,067.0. The variation suggests lexicon-specific perplexity spikes rather than semantic content differences.
  • Race. For race terms, black averages 6,945.6 while the archaic slur negroes averages 686.2, producing a 6,614.7 gap—evidence that token frequency and register strongly modulate the instrument.
  • Religion. Jewish terms show a 2,401.6 gap (jews vs. jew); Islamic vs. Arab terms differ by 457.8 in religion2.
Perplexity is sensitive to token rarity and spelling variants; nevertheless, such extreme disparities confirm annotation bias: the language model reacts to demographic tokens themselves, not merely to surrounding semantics. Consequently, any fairness evaluation that thresholds on perplexity (or that uses perplexity-based filters) will inherit these biases.
Researchers should therefore (i) Standardize demographic terms (e.g., map negroes\!$\to$ Black) before scoring; (ii) Calibrate or debias the LM on a demographically balanced corpus; (iii) Augment evaluation with human re-annotation to decouple token rarity from true offensiveness; (iv) Report all perplexity analyses alongside frequency statistics to avoid mis-attributing stylistic sparsity to social bias.
@@
Stereotype Leakage. We computed 9 496 (group, trait) co–occurrence pairs in the phrase split and visualised the results in Fig. 12.The analysis code will be released with the supplementary material. The overall mutual information is $\text{MI}=1.02$ bits—nearly an order of magnitude above a Wikipedia baseline—confirming that RedditBias is saturated with strong, non-neutral group–trait associations.
@@
  • Top stereotyped pairs. The bar-chart in Fig. 12 (upper left) confirms that mother–overprotective and mother–overbearing exhibit the highest leakage (PMI $\approx 6.45$), followed by overt Islamophobic and antisemitic pairings such as islamic–immigration and jewish–shooter.
@@
Differential Metric Bias. We examined score disparities between majority and minority subgroups across four demographic axes—gender, race, orientation, and religion—using pretrained metrics: Perspective API (toxicity family), VADER (sentiment), BERT-based Regard classifier, and a contextualized gender polarity scorer. These metrics were applied as defined in Equation (8), and results are visualized through boxplots and a differential heatmap.
We visualize three polarity measures in Figure 13: gender_unigram (based on word counts), gender_max (maximum contextual association with binary gender terms), and gender_wavg (weighted average of all gender associations in the sentence).
Gender Polarity. The unigram scores are nearly identical across all demographic groups, with values concentrated at discrete levels (e.g., 0, ±0.7). This is because the method simply counts gender-related keywords (e.g., “he,” “she,” “man”) and applies fixed polarity weights from a lexicon, independent of context, sentiment, or group identity. As a result, both minority and majority groups receive the same distribution of scores, rendering this metric ineffective for detecting differential metric bias.
In contrast, gender_max and gender_wavg are derived from contextualized language models and are sensitive to both word usage and framing. Minority subgroups—especially for gender and sexual orientation—exhibit higher maximum and average gender polarity scores, suggesting stronger or more stereotypically gendered language. This pattern supports the claim that bias in natural text is often embedded in subtle contextual cues, not surface-level term frequency.
Sentiment. Figure 14 shows that sentiment distributions are significantly more negative for race and orientation minorities. Sexual orientation minorities receive the lowest median and mean sentiment scores ($\Delta_{\text{sent}}=-0.054$), indicating that even short phrases involving those identities are evaluated by the VADER model as more negative. This pattern aligns with toxicity trends and suggests possible over-sensitivity of the sentiment metric to negatively valenced identity contexts.
Figure 15: Regard score distributions across demographic groups. In contrast to sentiment and toxicity, regard scores are relatively invariant between majority and minority groups.
Regard. Despite variations in sentiment and toxicity, regard scores remain relatively stable across all groups, as seen in Figure 15. The maximum groupwise difference is less than 0.02 points, and the distributions of majority and minority subgroups are nearly overlapping. This suggests that the BERT-based regard classifier is less sensitive to lexical signals of identity and may not reflect implicit polarity or connotation differences in conversational data. While this stability is favorable in terms of fairness, it may also indicate a lack of sensitivity to harmful nuance.
Toxicity and Related Metrics. Perspective API scores show the most pronounced differential bias across axes, especially for sexual orientation and race. Figure 16 illustrates that orientation minorities receive significantly higher scores across all toxicity dimensions, with mean differences up to +0.125 in general toxicity and +0.146 in identity attack. Race minorities also exhibit elevated levels in severe toxicity and insult. Notably, gender toxicity differences are minimal. These findings are corroborated in the summary heatmap (Figure 17), where all values are negative (majority - minority), confirming that minority groups consistently receive higher toxicity-related scores.
Our analysis confirms that RedditBias exhibits strong differential metric bias across several demographic axes. Toxicity and identity-related metrics show the most severe disparities, especially for orientation and race. Sentiment scores are also directionally aligned with these results, while regard remains relatively stable. Gender polarity measures indicate that minority group references are more gendered in context. These findings suggest that evaluation pipelines using these metrics should account for metric-induced bias to avoid amplifying false disparities, especially in safety-critical dialogue systems.

### v2

@@ 3.3.2 CrowS-Pairs
Dataset Taxonomy. CrowS-Pairs is classified as a constrained-form evaluation dataset within the sentence-likelihood counterfactuals category. It uses contrastive sentence pairs where the only substantive difference between each pair is the demographic group mentioned, with one sentence referencing a historically marginalized group and the other referencing a contrasting advantaged group. The dataset evaluates whether language models systematically favor stereotypical sentences by assigning them higher likelihood, with an unbiased model expected to assign equal or lower likelihood to the stereotypical version, reflecting social fairness principles.
In terms of source, CrowS-Pairs employs crowdsourced construction through Amazon Mechanical Turk. Workers were provided with partial prompts from sources like MultiNLI or ROCStories and instructed to compose minimal pairs reflecting societal biases. Substitutions were guided by a predefined list of group pairs, and all submissions underwent a rigorous validation process involving five annotators. This approach incorporates diverse viewpoints from annotators while potentially introducing inconsistencies in quality and demographic biases from the annotators themselves.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on American English with stylistically diverse content including narrative, conversational, and factual tones. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, CrowS-Pairs targets multiple demographic characteristic biases across nine categories: race/color, gender/gender identity, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each pair is tagged with a primary bias type, providing comprehensive coverage of protected attributes. However, the dataset’s construction may introduce dataset construction biases, particularly annotation bias from subjective judgments of crowdworkers and potential stereotype leakage in the framing of sentence pairs.
Finally, CrowS-Pairs demonstrates high accessibility as a publicly available dataset. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Representativeness Bias. Following the methodology defined in Section 2.3, we assess representativeness bias in CrowS-Pairs by comparing its demographic distribution to U.S. population priors from the 2020 Census (for race and gender), 2022 American Community Survey (for nationality and disability), Pew Research Center 2021 (for religion and income), Gallup 2023 (for sexual orientation), and CDC 2021 BMI statistics (for weight status). Let $P_{\mathcal{D}}(A)$ denote the empirical distribution across demographic categories and $P_{\mathcal{P}}(A)$ denote the reference distribution. The representativeness bias is computed as Equation 1, the Kullback–Leibler divergence between these two distributions. This metric reflects the extent to which the dataset over- or under-represents certain demographic groups relative to real-world proportions.
i) Race–Color. CrowS–Pairs significantly over-samples sentences targeting Black individuals (40.1% vs. 13.6%), while under-sampling White (44.3% vs. 58.9%) and Hispanic groups (5.3% vs. 19.1%). This results in a high divergence of
@@
ii) Gender. The gender axis is more balanced, with a slight male overrepresentation (55.1% vs. 51.2%) and female underrepresentation (44.9% vs. 48.8%), yielding
@@
iii) Other Attributes. While CrowS–Pairs lacks structured subgroup labels for attributes like socioeconomic status, nationality, disability, and religion, visual comparison with the BLS 2023 reference distributions (see Figure 10) reveals similar mismatches:
@@
For detailed demographic statistics used to derive these reference values, see Appendix A.1.2.
@@
Figure 10: CrowS–Pairs (blue) vs. 2023 BLS demographic distributions (red) across nine protected attributes. Over- and under-representation are evident where bars differ substantially.
Annotation Bias. CrowS–Pairs depends on crowdworkers to create and label each contrastive pair, designating one sentence as stereo (stereotype) and the other as anti_stereo (counter-stereotype). Using the annotation-bias estimator from Equation (2), we audited all 1,508 validated pairs and obtained two main findings.
To probe whether annotators systematically treat certain demographic categories as more “stereotype‑prone” than others, we begin by auditing the distribution of stereotypical labels across and within attributes.
  • Across bias types. Stereotypical‐label prevalence varies sharply. Disability (0.9500), religion (0.9429), and race-color (0.9167) are most frequently tagged as stereotypical, whereas gender (0.6069) and physical appearance (0.8254) are lower. The maximum inter-type gap is
@@
    occurring between gender and disability.
  • Within attributes. Inside race-color the rates are:
@@
    The intra-attribute gap is $B_{\mathrm{ann}}^{\text{group}}=0.0816$, between Asian and Other. For gender, stereotypical labels are 0.8440 for male vs. 0.8031 for female.
These disparities show that annotators tend to flag sentences about disability, religion, and race‐color as stereotypical far more often than those about gender, indicating unequal sensitivity across social dimensions.
Complementing the gold‑label audit, we next examine how external scoring tools treat the same pairs. We computed the paired score gap $\Delta_{f}=f(x_{\text{stereo}})-f(x_{\text{anti}})$ for seven popular metrics over all pairs and averaged them by bias type. Figure 11 visualises the results.
Figure 11: Mean paired score gaps $\mathbb{E}[\Delta_{f}\mid\text{bias\_type}]$. Blue (red) denotes lower (higher) scores for stereotype sentences.
After Bonferroni correction (p>0.05/7p{>}0.05/7), none of the Perspective-API facets, VADER sentiment, or Regard shows a significant global gap,  [^7] Toxicity: $\bar{\Delta}=0.021$, $t=1.37$, $p=0.18$; Sentiment: $\bar{\Delta}=-0.039$, $t=0.04$, $p=0.97$; Regard: $\bar{\Delta}=0.0004$, t=1.49t=-1.49, $p=0.14$. yet the embedding-based gender-polarity suite yields small but reliable negative shifts (unigram d=0.26d=-0.26, max d=0.34d=-0.34, wavg d=0.28d=-0.28, all $p<5{\times}10^{-4}$), meaning stereotype sentences lean slightly masculine in embedding space.
Figure 11 reveals that metric bias is heterogeneous across bias types:
  • Gender rows illustrate the massive negative gender_unigram gap (0.261-0.261), corroborating the aggregate significance of gender polarity.
Viewed together, the dual biases uncovered above motivate several precautions for researchers who rely on CrowS‑Pairs in fairness studies: (1) gold-label skew that over-stigmatizes certain categories, and (2) non-uniform reactions from common scoring instruments. Hence, fairness evaluations that depend on this dataset should (1) report both global and per-category statistics, (2) audit the auxiliary metrics themselves, and (3) interpret gender findings in light of the metrics’ masculine tilt.
Stereotype Leakage. We computed pointwise mutual information (PMI) for every (group, trait) co-occurrence using a fixed vocabulary of protected-group and trait tokens, defined in Appendix A.2.888The word lists span five protected categories (race–color, gender, religion, age, and disability) and four trait domains (occupation, personality, social_status, and behavior). See Appendix B.3 for full listings and implementation notes. Figure 12 visualizes the results.
@@
  • Mutual-information ranking (bits): gender–personality ($1.04$) >> gender–behavior ($0.92$) >> race/social_status ($0.54$) >> religion/occupation ($0.50$); combinations such as race–behavior and age–social_status yield $\mathrm{MI}=0$, reflecting either scarcity or filtering.
@@ 3.3.3 RedditBias
Dataset Taxonomy. RedditBias is classified as a constrained-form evaluation dataset within the sentence-likelihood counterfactuals category. Unlike synthetic datasets, it captures naturally occurring biased language from real-world Reddit discussions, providing two levels of granularity: full Reddit comments and phrase-level windows around target demographic terms. The dataset focuses on overt, toxic stereotypes and avoids subtle bias expressions, making it particularly suited for evaluating how dialogue models handle harmful content in conversational settings while maintaining ecological validity through authentic user-generated text.
In terms of source, RedditBias employs natural text sources, compiled using 3.33 years of Reddit data obtained via the Pushshift API. This approach provides rich linguistic diversity and represents real-world language usage, offering a more comprehensive view of how biases might manifest in actual communication. However, the challenge with natural text sources is that they may already contain societal biases, making it difficult to discern whether bias comes from the model itself or the underlying data.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on informal American English with content rich in sarcasm and slurs, reflecting the authentic linguistic style of Reddit. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, RedditBias targets multiple demographic characteristic biases across four axes: Gender, Race, Religion, and Queerness. The dataset focuses on overt, toxic stereotypes and avoids subtle or positive bias expressions, using curated lists of group identifiers and negative attributes. However, the dataset’s construction may introduce dataset construction biases, particularly representative bias from non-representative sampling of Reddit content and potential stereotype leakage in the framing of retrieved phrases.
Finally, RedditBias demonstrates restricted accessibility, distributed under a CC BY-NC-SA license and including verbatim Reddit content, which may limit certain applications. This restricted access contrasts with fully public datasets but still enables research use while addressing potential privacy and content concerns.
@@
Strengths and Limitations. A key strength of RedditBias lies in its realism, as it reflects how bias and hate speech appear in natural online dialogue. The dual-level design (comment and phrase) enables both coarse and fine-grained assessments. It also supports both intrinsic scoring using model-likelihood metrics and downstream evaluations of dialogue performance. However, the dataset has several limitations. Its domain specificity may reduce generalizability to other platforms or languages. It focuses solely on four demographic dimensions and explicitly targets harmful bias, omitting subtle or structural forms. The reliance on pattern-matching during data collection may overlook implicit stereotypes, and the use of sarcasm or ambiguous phrasing complicates annotation and model interpretation. Additionally, the dataset contains highly offensive content, which necessitates ethical caution in use and publication. Its composition is heavily skewed toward toxic expressions, without balancing benign or neutral dialogue, which may influence evaluation metrics.
Bias Analysis. Representativeness Bias. Using U.S. Census 2020 (for race and gender), 2022 American Community Survey (for nationality and disability), Pew Research Center 2021 (for religion and income), Gallup 2023 (for sexual orientation), and CDC 2021 BMI statistics (for weight status) as reference benchmarks, we recomputed the KL divergence between the RedditBias attribute distribution and the reference population, obtaining
$B_{\text{rep}}^{\text{RedditBias}}=D_{\mathrm{KL}}\!\bigl(P_{\mathcal{D}}(A)\,\|\,P_{\text{BLS 2023}}(A)\bigr)=17.75,$ (31)
far larger than both our earlier estimate (14.51) and the Wikipedia‐based baseline reported in this section. The principal drivers of this divergence are:
  • Gender. Female‐coded terms (girl, mom, wife, lady) dominate, whereas BLS 2023 reports a near‐parity labour distribution (51.2 % male, 48.8 % female).
  • Sexual Orientation. LGBTQ + terms (gay, transgender, lesbian) appear in $\approx$ 7× the proportion suggested by BLS 2023 figures (7.2 % LGBTQ +, 92.8 % straight).
  • Race / Ethnicity. Black-related tokens are heavily over-sampled; Whites (58.9 %) and Hispanics (19.1 %) are markedly under-represented relative to BLS 2023 population shares.
  • Religion. Jewish and Islamic terms occur much more often than their combined 37 % share of non-Christian affiliations (BLS 2023: 63 % Christian, 37 % non-Christian).
@@
Annotation Bias. Under our unified definition (Section 2.3), annotation bias arises whenever the labeling function $g_{\theta}(x)$, be it manual or automatic, assigns systematically different labels to semantically comparable instances across demographic groups. In RedditBias, we operationalize $g_{\theta}$ with several automatic scorers that are frequently used as proxy labels or filters in evaluation pipelines: (i) sentence-level perplexity from a 6-layer GPT-2 fine-tuned on balanced news, (ii) the Perspective API toxicity family, (iii) VADER sentiment, (iv) a BERT-based regard classifier, and (v) three gender-polarity metrics (one lexicon-based, two contextual). Because these scorers can be, and often are, used to label or threshold instances, any systematic disparity they introduce is, by our definition, a form of annotation bias.
We first instantiate $g_{\theta}(x)=\mathrm{ppl}(x)$ and estimate Equation (2). The maximum groupwise gap is
@@
which far exceeds the 95% bootstrap confidence band for neutral vocabulary ($\approx 1200$). Across axes, we observe extreme and often unintuitive disparities that reflect the scorer’s sensitivity to token rarity, register, and spelling variants rather than to semantics alone. For instance, within gender, the largest gap (23,675.67) occurs between niece and bride; class means span from 803.4 (niece) to 8,424.4 (mother), indicating that the LM treats female-coded tokens as substantially harder to predict. In orientation, homosexuals and bisexuals differ by 3,622.6, with lesbian at 4,067.0, suggesting lexicon-specific spikes. For race, black averages 6,945.6 whereas the archaic negroes averages 686.2, producing a 6,614.7 gap, evidence that frequency and register strongly modulate perplexity. Jewish terms show a 2,401.6 gap (jews vs. jew); Islamic vs. Arab terms differ by 457.8 in religion2. These results demonstrate that perplexity-based filters or thresholds will systematically over-select or under-select content associated with particular identity markers, thereby injecting annotation bias into any downstream dataset that relies on such scoring.
Turning to other automatic scorers, Figures 1315 document substantial cross-group disparities. The lexicon-based gender_unigram polarity metric yields nearly identical, discretized distributions across groups (e.g., $0,\pm 0.7$}), revealing its insensitivity to context and, thus, its limited diagnostic value for annotation bias. In contrast, the contextualized gender_max and gender_wavg scores capture stronger gender associations for minority-marked text (Figure 13), implying that these labeling functions encode contextual and framing effects that differ by group.
Sentiment (Figure 14) is systematically more negative for orientation and race minorities, with sexual-orientation minorities receiving the lowest median and mean scores ($\Delta_{\text{sent}}=-0.054$). This indicates that even short phrases involving these identities are assigned more negative polarity by VADER, introducing group-dependent label skew if sentiment is used as a proxy outcome. By contrast, regard scores are relatively invariant, with a maximum groupwise gap below 0.02, which is favorable in terms of fairness, but possibly under-sensitive to harmful nuance in conversational settings.
The strongest disparities arise in the Perspective API family (Figure 16), where orientation and race minorities receive consistently higher scores across toxicity, severe toxicity, identity attack, insult, and threat. Mean differences reach up to +0.125 (toxicity) and +0.146 (identity attack), respectively, and the heatmap in Figure 15 shows uniformly negative majority–minority deltas, confirming that minority groups are systematically scored as more toxic. When these toxicity scores are used to label, filter, or weight examples, they induce precisely the kind of cross-group label disparity captured by Equation (2).
Taken together, these findings show that RedditBias, when paired with widely used automatic scorers, exhibits pronounced annotation bias: the labeling functions $g_{\theta}$ (perplexity, toxicity, sentiment, contextual polarity) are themselves demographically sensitive. To mitigate this, we recommend: (i) standardizing and consolidating demographic lexical items (e.g., mapping negroes $\to$ Black) before scoring; (ii) calibrating or debiasing scorers on demographically balanced corpora; (iii) supplementing automatic scores with targeted human re-annotation to disentangle token frequency from genuine offensiveness or harm; and (iv) always reporting frequency statistics and confidence bands alongside groupwise score gaps to avoid attributing sparsity- or register-driven effects to social bias. Without such controls, any evaluation pipeline that thresholds or optimizes on these automatic outputs risks hard-coding annotation bias into its conclusions.
@@
Stereotype Leakage. We computed 9 496 (group, trait) co–occurrence pairs in the phrase split and visualised the results in Figure 17.The overall mutual information is $\text{MI}=1.02$ bits—nearly an order of magnitude above a Wikipedia baseline—confirming that RedditBias is saturated with strong, non-neutral group–trait associations.
@@
  • Top stereotyped pairs. The bar-chart in Figure 17 (upper left) confirms that mother–overprotective and mother–overbearing exhibit the highest leakage (PMI $\approx 6.45$), followed by overt Islamophobic and antisemitic pairings such as islamic–immigration and jewish–shooter.